A mother is presently four times as old as her daughter.If the sum of their ages 6 years ago was 48, find their present ages
Answer by stanbon(75887)

A mother is presently four times as old as her daughter.If the sum of their ages 6 years ago was 48, find their present ages
-------
Equations:
m = 4d
m-6 + d-6 = 48
--------------------
Substitute for "m" and solve for "d":
4d-6 + d-6 = 48
5d - 12 = 48
5d = 60
d = 12 years (daughter's age now)
------
Solve for "m":
m = 4d
m = 4*12 = 48 (mom's age now)
===========
Cheers,
Stan H.
=============